If 5.23 g of CuNO3 is dissolved in water to make a 0.800 M solution, what is the volume of the solution in milliliters?

volume
Answer #1

molarity = number of moles / volume of solution in L

number of moles = 5.23 g / 125.5509 g/mol = 0.0417 mole

molarity = number of moles / volume of solution in L

0.800 = 0.0417 / volume of solution in L

volume of solution in L = 0.0417 / 0.800 = 0.0521 L

1 L = 1000 mL

0.0521 L = 52.1 mL

Therefore, the volume of solution = 52.1 mL